### v1

# 1]organization=Scuola Normale Superiore and INFN, addressline=Piazza dei Cavalieri 7, city=Pisa, postcode=56126, state=PI, country=Italy

Abstract
We explore the phenomenology of the “Dark Walker”—an $\text{SU}(3)$ theory with eight flavors of fundamental fermions in the dark sector. During inflation, its walking dynamics generate primordial non-Gaussianities through the exchange of unparticles, while accounting for the current dark matter relic abundance if the resulting massive pseudoscalars behave as strongly interacting massive particles (SIMPs). This provides a simple yet concrete example of a strongly coupled sector during inflation.
keywords:
[…]
Introduction.—Assuming that the early universe underwent a phase of cosmic inflation, fields that weakly couple to the inflaton can play a crucial role in shaping the primordial non-Gaussianities—also referred to as cosmological correlators. One possible origin of scalar perturbations during inflation is the presence of composite particles, which arise from strongly coupled dynamics among more elementary constituents, similar to the mechanisms in quantum chromodynamics (QCD) and models of the composite Higgs. This scenario has been far less explored compared to the more conventional treatments of cosmological correlators involving fundamental particle exchanges. For recent developments in this direction, see, for example, [1, 2, 3, 4, 5].
In this Letter, we consider a dark sector that interacts with the Standard Model (SM) through the Higgs portal and becomes strongly coupled during inflation. After inflation, this dark sector contributes to today’s relic abundance of dark matter (DM). As a simple illustrative example, we study what we call the “Dark Walker”—a scenario that has the potential to account simultaneously for non-Gaussianities and the observed DM relic abundance.
Setup.—To investigate a strongly coupled dark sector during inflation, it is useful to classify different scenarios according to the mass gap: for example, $\text{M}_{\text{Gap}}\ll H_{\text{inf}}$, $\text{M}_{\text{Gap}}\sim H_{\text{inf}}$ and $\text{M}_{\text{Gap}}\gg H_{\text{inf}}$. In this work, the Dark Walker model falls into the $\text{M}_{\text{Gap}}\ll H_{\text{inf}}$ regime, where the dark sector is well approximated by a conformal field theory. Such a conformal sector may contain operators with large anomalous dimensions, which are called unparticles, as introduced in [6, 7], etc. An extended notion of unparticles with a mass gap has also been discussed in [8], but lies beyond the scope of this work. Gapless unparticles, however, cannot serve as dark matter, since their behavior resembles radiation rather than particles. Thus, the Dark Walker has to experience a phase transition, forming bound states. During inflation, the model remains near the infrared fixed point (IRFP) long enough to leave imprints in the primordial non-Gaussianity. This behavior exemplifies a form of “walking dynamics”, which gives rise to the name “Dark Walker”.
Let us consider a non-Abelian gauge theory in the dark sector (DS), which is an $\text{SU}(N_{c})_{D}$ theory with $N_{f}$ flavors of massless Dirac fermions. At the start of inflation, all modes are very blue-shifted and the effect of the de Sitter curvature in the $\beta$-function will be small, so we can approximately treat the background as flat space. The Lagrangian is
[…]
The $T_{a}$’s are the matrices of the adjoint representation of $\text{SU}(N_{c})_{D}$ and $g$ is the gauge coupling.
It is long known that there exists a non-trivial IRFP for $\text{SU}(N_{c})$ theories with a certain range of $N_{f}$, which is called the conformal window [9]. The $\beta$-function of $\text{SU}(N_{c})$ theories in the asymptotically free phase is given in [10, 11]
$\beta(g)=-\beta_{0}\frac{g^{3}}{(4\pi)^{2}}+\beta_{1}\frac{g^{5}}{(4\pi)^{4}}+O(g^{7}),$ (2)
where $\beta_{0}$ and $\beta_{1}$ are
$\beta_{0}$ $=\frac{11}{3}N_{c}-\frac{2}{3}N_{f},$ (3)
$\beta_{1}$ $=-\frac{34}{3}N_{c}^{2}+\frac{10}{3}N_{c}N_{f}+\frac{N_{c}^{2}-1}{N_{c}}N_{f}.$ (4)
The theory could arrive at an IRFP if the first two terms have opposite signs. Up to two loops, the coupling constant $g_{c}^{2}$ will be
$\frac{g_{c}^{2}}{(4\pi)^{2}}=\frac{\beta_{0}}{\beta_{1}}=\frac{11N_{c}^{2}-2N_{c}N_{f}}{13N_{c}^{2}N_{f}-3N_{f}-34N_{c}^{3}}.$ (5)
From the $\beta$-function we notice that, generally in the perturbative regime $g\lesssim 4\pi$, the IRFP is stable [9], which means that a small relevant deformation will not drag the Dark Walker into the gapped phase. Therefore, we have to look for special gauge theories. Fortunately, recent studies in lattice QCD (e.g. [12, 13]) have revealed that, for the SU(2) theory with four flavors of massless Dirac fermions and the $\text{SU}(3)$ theory with eight flavors of massless Dirac fermions, the ultraviolet (UV) and IR fixed points will merge, as illustrated in Figure 1.
Figure 1: The $\beta$-function with merged fixed points
In the following, we will take the Dark Walker to be an $\text{SU}(3)_{D}$ theory with $N_{f}=8$ flavors. This is an illustrator of our Dark Walker mechanism. In general, this scenario should hold for all gauge models with MFP.
Into inflation.—In this work, we choose the value of $H_{\text{inf}}$ to be one power of ten times smaller than the upper bound value from [14], which is $H_{\text{inf}}\sim 10^{-1}\times 2.5\times 10^{-5}M_{\text{pl}}\sim 6.0\times 10^{12}\text{ GeV}$. In the context of cosmology, it is convenient to use the number of $e$-folds $N$ to measure the ratio between scales. There are $\sim 13$ $e$-folds between the Planck scale and the Hubble scale at the start of inflation.
If we start from a weak gauge coupling in the UV, we can estimate the number of $e$-folds needed to arrive at the merged fixed point (MFP) using
$\frac{\text{d}g}{\text{d}N}=\beta_{0}\frac{g^{3}}{(4\pi)^{2}}-\beta_{1}\frac{g^{5}}{(4\pi)^{4}}$
$\Rightarrow\ N_{\text{cw}}&=\frac{8\pi^{2}}{\beta_{0}}\left(\frac{1}{g_{0}^{2}}-\frac{1}{g_{c}^{2}}\right)\\ &+\frac{\beta_{1}}{2\beta_{0}^{2}}\log\left(\frac{g_{c}^{2}}{g_{0}^{2}}\frac{16\pi^{2}\beta_{0}-g_{0}^{2}\beta_{1}}{16\pi^{2}\beta_{0}-g_{c}^{2}\beta_{1}}\right).\end{aligned}$ (6)
We will take $g_{c}^{2}$ to be the lattice result of [13], where $g_{\text{GF}}^{2}\approx 25$ will give us $g_{c}^{2}\lesssim 12$, with reference to [15]. Surely, this one-loop matching in [15] will break up for strong couplings, but all we need is an upper bound for the number of $e$-folds needed.
Since $g_{c}/4\pi\lesssim 1$, (6) is enough for an estimate. Based on the number of $e$-folds between $M_{\text{pl}}$ and $H_{\text{inf}}$, to prepare the unparticles, we expect less than 10 $e$-folds to achieve the MFP. Taking $g_{0}=2$, the number of $e$-folds needed and the resulting energy scale for forming the conformal window will be
$&N_{\text{cw}}\approx 4.0,\\ \Rightarrow\ &\Lambda_{\text{cw}}=e^{-4.0}M_{\text{pl}}\approx 4.4\times 10^{16}\text{ GeV}.\end{aligned}$ (7)
Once our model reaches the MFP, the gauge theory (1) becomes a conformal theory, and we cannot use this Lagrangian. In general, it is not easy to derive the operator spectrum and the operator product expansion (OPE) coefficients between operators from the UV theory (1). Nevertheless, we can focus on a few operators that we expect to see from the UV theory. An important one is the primary scalar operator $\mathcal{O}_{\psi}\equiv\bar{\psi}\psi$, as the anomalous dimension $\gamma_{m}$ of $\mathcal{O}_{\psi}$ will be greater than 1 at the MFP, studied in [16]. This property will trigger “walking” and drag the Dark Walker into the gapped phase, as we will see below. Another possible primary scalar operator is $\mathcal{O}_{F}\equiv\text{Tr}F^{2}$, where the scaling dimension $\Delta_{F}\lesssim 4$. The explicit value can only be found from lattice simulations. The operator $\mathcal{O}_{\psi}$ has the scaling dimension $\Delta_{\psi}=3-\gamma_{m}\approx 1.9$ from [16], and we assume $\Delta_{F}\approx 3.9$ for $\mathcal{O}_{F}$.
Using the language of the effective field theory (EFT) of inflation [17, 18], we can construct the interactions between the scalar primary operators $\mathcal{O}_{\Delta}$ and the scalar perturbation $\pi$ as given in [1, 5]:
$\mathcal{L}^{(s=0)}_{\pi\mathcal{O}}\equiv\frac{1}{2}\mu^{2-\Delta}M_{\text{pl}}|{\dot{H}_{\text{inf}}}|^{1/2}\mathcal{O}_{\Delta}(g^{00}+1)$
$=\frac{1}{2}\mu^{2-\Delta}M_{\text{pl}}|{\dot{H}_{\text{inf}}}|^{1/2}\left(-2\dot{\pi}\mathcal{O}_{\Delta}+(\partial_{\mu}\pi)^{2}\mathcal{O}_{\Delta}\right).$ (8)
It is worth mentioning that now we should regard the unparticles $\mathcal{O}_{\psi}$ and $\mathcal{O}_{F}$ as composite particles, which means, for example, $\dot{\pi}\mathcal{O}_{\psi}$ is a linear mixing instead of a Yukawa coupling. The dimension of the linear mixing $\dot{\pi}\mathcal{O}_{\psi}$ is smaller than four, making this linear mixing relevant. This interaction will walk the theory from the MFP into the gapped phase, while a small relevant interaction will not affect the overall dynamics of inflation.
Since now the power of the coupling constant depends on $\Delta_{\psi}$ and $\Delta_{F}$, we need the effective coupling constant $\mu_{\psi}$ to be much smaller than $H_{\text{inf}}$ and $\mu_{F}$ to be much larger than $H_{\text{inf}}$, in order to have a sufficient perturbative description. In a slow-roll model, we can take
$\mu_{\psi}$ $=|\dot{H}_{\text{inf}}|/\Lambda_{\text{cw}}\approx 7.9\times 10^{6}\text{ GeV}\ll H_{\text{inf}},$ (9)
$\mu_{F}$ $=\dot{\phi}/\Lambda_{\text{cw}}\approx 4.6\times 10^{13}\text{ GeV}\gg H_{\text{inf}},$ (10)
which is good enough for perturbative expansion. In addition, assuming that $\langle\mathcal{O}_{\Delta}^{3}\rangle$ is negligible, the bispectrum is determined by the quadratic derivative coupling $(\partial_{\mu}\pi)^{2}\mathcal{O}_{\Delta}$, as discussed in [5].
Non-Gaussianity.—The cosmological bispectrum can be calculated using the results in [5]. The non-Gaussianity $f_{\text{NL}}$ is defined as
$f_{\text{NL}}\equiv\frac{5}{18}\frac{\mathcal{B}_{\zeta}(k,k,k)}{\Delta_{\zeta}^{4}(k)},$ (11)
where $\mathcal{B}_{\zeta}$ is the bispectrum of curvature perturbation $\zeta$, and $\Delta_{\zeta}^{2}=2.1\times 10^{-9}$ is the dimensionless power spectrum of $\zeta$ from [19]. The $f_{\text{NL}}$ using the language of the EFT of inflation can be written as
$f_{\text{NL}}=\frac{5}{18}\frac{1}{\Delta_{\zeta}^{4}}\left(\frac{1}{2}\mu^{2-\Delta}M_{\text{pl}}|{\dot{H}_{\text{inf}}}|^{1/2}\right)^{2}H_{\text{inf}}^{2\Delta-1}$
×$\times\frac{1}{M_{\text{pl}}^{2}|{\dot{H}_{\text{inf}}}|}\left(-\frac{H_{\text{inf}}}{\dot{\phi}}\right)^{3}\times\frac{(2\pi)^{4}}{16k^{6}}S(k,k,k).$ (12)
Taking the slow-roll parameter $\epsilon$ to be $0.0097$ according to [14], we find that $f_{\text{NL},\psi}\approx 16$ and $f_{\text{NL},F}\approx 3.5\times 10^{-3}$, which are desirable because they are around the center of the latest observation $f_{\text{NL}}=6\pm 46$, as shown in [20]. Notice that the choice of $H_{\text{inf}}$ will significantly affect the value of $f_{\text{NL}}$. Figure 2 shows the shape functions with $\mathcal{O}_{\psi}$ and $\mathcal{O}_{F}$ exchanges. Both of them are near-orthogonal, while the squeezed limit behaviors degenerate with the equilateral non-Gaussianity. Only the full shapes of the bispectra can break this degeneracy, as emphasized in [5].
Figure 2: Shape functions for $\Delta=1.9$ and $\Delta=3.9$. We rescale the magnitudes by $10^{2}$ for $\Delta=1.9$ and $10^{3}$ for $\Delta=3.9$ to set the plot value within order 1.
After inflation.—After the inflation ends, the universe becomes almost flat. The interactions $\dot{\pi}\mathcal{O}_{\psi}$ and $(\partial_{\mu}\pi)^{2}\mathcal{O}_{\psi}$ will freeze, since the energy scale becomes too small compared to the EFT of inflation scale $M_{\text{pl}}|{\dot{H}_{\text{inf}}}|^{1/2}$. However, even though the linear mixing interaction $\dot{\pi}\mathcal{O}_{\psi}$ remains small before it gets frozen, its nature as a relevant operator will drag our model into the gapped phase, creating bound states. They are natural dark matter candidates. We expect them to be standard SIMP-like dark matter. When we take the ratio between the dark matter mass and the freeze-out temperature to be the standard choice $x_{f}\equiv m_{\text{DM}}/T_{f}\sim 20$, we can believe that the lightest state dominates. By taking the freeze-out temperature to be below the QCD phase transition scale, the dark sector will remain in thermal equilibrium with the SM particles before freeze-out. The masses of the lowest lying states of pseudoscalars (PSs) and vector mesons (Vs) can be read from [13, 21]. The mass of the lightest vector mesons is 2.2 times the mass $m_{\text{PS}}$ of the lightest pseudoscalars, which means that the vector mesons are not the light vector mesons discussed in [22]. The pseudoscalars will dominate the relic abundance, since the vector mesons freeze out first. We can infer the mass scale of the dark pseudoscalars from the observed relic abundance. Since the bound states come from a small relevant deformation-induced confinement, the interactions in the conformal phase have negligible effects on the properties of bound states.
Based on the standard SIMP discussions in [23, 24], the dominant annihilation cross section is the $3\rightarrow 2$ self-annihilation cross section $\langle\sigma v^{2}\rangle_{3\rightarrow 2}$. It can be parametrized by an effective coupling $\alpha_{\text{eff}}$ in the way
$\langle\sigma v^{2}\rangle_{3\rightarrow 2}\equiv\alpha_{\text{eff}}^{3}/m_{\text{PS}}^{5}.$ (13)
Taking the DM to be non-relativistic at the freeze-out, the effective number of degrees of freedom is $g_{*}=10.75$.
[25] gives the expression for computing the relic abundance when the $3\rightarrow 2$ annihilation dominates, as
$\Omega_{\text{DM}}h^{2}=\frac{1.1\times 10^{-10}\text{ GeV}^{-2}}{\left(g_{*}^{3/2}\frac{m_{\text{PS}}^{2}}{M_{\text{pl}}}\int_{x_{f}}^{\infty}\frac{\text{d}x}{x^{5}}\left\langle\sigma v^{2}\right\rangle_{3\rightarrow 2}\right)^{1/2}}.$ (14)
To have the present-day observed value $\Omega_{\text{DM}}h^{2}=0.12$ [19], we found
$m_{\text{PS}}\sim\alpha_{\text{eff}}\times 30\text{ MeV}.$ (15)
If we take $\alpha_{\text{eff}}=1$ as in [23], $m_{\text{PS}}$ becomes $\sim 30\text{ MeV}$, which is a proper mass scale for SIMP dark matter.
Discussion.—A simple gauge theory at the lower edge of the conformal window in the dark sector opens up new possibilities for beyond-Standard-Model (BSM) physics and dark matter candidates. The Dark Walker model has the potential to produce both the primordial non-Gaussianity, which is large enough to have the possibility of being detectable, and the observed relic abundance of dark matter. As expected, several parameters must still be fixed by hand, though their values remain within natural ranges. The non-Gaussianity signal is sensitive to the inflationary Hubble scale, the initial gauge coupling—which determines the starting energy scale of the merged fixed point—and the precise anomalous dimensions of the primary scalar operators in the conformal phase. These parameters may be constrained more accurately in the future through improved observations of inflationary parameters and lattice QCD studies of the spectrum in the $\text{SU}(3)$ theory with eight flavors. With such inputs, more precise predictions of the primordial non-Gaussianity and relic abundance can be achieved. This framework offers a concrete testing ground for the role of strongly coupled sectors in primordial cosmology.
Acknowledgements
I express my thanks to Riccardo Barbieri, who encouraged me very helpfully while this work was in progress. Thanks to Raffaele Tito D’Agnolo, Janet Hung, Guilherme Leite Pimentel, and Alessandro Strumia for useful discussions and comments.
CY is supported by Scuola Normale and by INFN (IS GSS-Pi). The research of CY is moreover supported by the ERC (NOTIMEFORCOSMO, 101126304). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.